•Susan put $375 into an account and one year later had $405. What interest rate was paid on Susan’s deposit?

•A. 5 percent

•B. 7 percent

•C. 8 percent

•D. 10 percent
Answer #1

Answer : Option C is​​​​​ correct.

A=$405

P =$375

n= 1 year

A= P(1+r)n

405= 375( 1+r)1

1.08= 1+r

r=0.08

Therefore, rate of interest is 8%